Hi,
I would really appreciate some help with this question:
A painter can prepare and paint a room in 10 hours. His son requires 18 hours to do the same job. How long will it take to prepare and paint the room working together?
Thank you,
Louis

Answer by Earlsdon(6294)

Here's one approach: Find the hourly rates of each person.
If the painter can do the whole job in 10 hours, then he can do 1/10 of the job in 1 hour.
If his son can do the whole job in 18 hours, then he can do 1/18 of the job in 1 hour.
Together, they can do (1/10 + 1/18) of the job in 1 hour.
So together, in 1 hour they can do (9/90 + 5/90 = 14/90) of the job in 1 hour.
This means it will take them 90/14 hour to do the whole job together.
90/14 hour = 6 hrs and 25.7 minutes.